### v1

@@ Abstract
Magnetic reconnection in magnetized wakes of cosmic strings results in the release of a large amount of energy. This energy is released in a short period of time. In this work, we show that this sudden release of energy can result in a Gamma Ray Burst (GRB) of short duration. The magnetic reconnection occurs at several points of the cosmic string wake. The emerging shocks from these points have different velocities. These shocks will collide with each other and give rise to short bursts of energy. The emitted pulse of energy depends on the background magnetic field and the timescale associated with the magnetic reconnection in the cosmic string wake. We also obtain the synthetic lightcurve that occurs from multiple collisions of the shock waves that are emitted from the multiple points of magnetic reconnection in the wake region. Finally, we discuss about some current experimental evidence of short GRB’s which are in the same energy range as the ones predicted from our model.
@@ I Introduction
Symmetry breaking phase transitions in the early universe give rise to topological defects. One of the topological defects that is predicted by these symmetry breaking phase transitions is the cosmic string defect. The cosmic string defect is a one dimensional defect generated due to a $U(1)$ symmetry breaking phase transition. Once formed, these strings will move through the early universe plasma. Depending on the exact nature of the phase transition, long strings or loops of cosmic strings are formed. Some of these defects are unstable and decay with the evolution of the universe, others are stable and evolve as a giant network of cosmic strings spanning the universe cosmicstrings . Theoreticians have predicted multiple observational signatures of these cosmic strings. Primary amongst these were the gravitational lensing of objects due to the conical space time of the cosmic string lensing . Other than the gravitational lensing, as the cosmic string moves through the plasma, it gives rise to wakes of overdensity behind it wakes . These wakes will lead to a distinct wedge shaped patterns in the 21 cm redshift survey maps branden1 ; dacunha . These wakes also give rise to temperature fluctuations which have been used for searches for these exotic objects by the Planck collaboration Planck_2016 . All these observational signatures are currently from density inhomogeneities generated as the cosmic string moves through the plasma.
@@
The wakes of these cosmic strings are long and narrow. The Biermann battery mechanism ohira produces magnetic fields that are oppositely directed along the two sides of the wake. A detailed study using the Sweet-Parker model has shown the possibility of magnetic reconnection in these magnetized string wakes Dilip_2023 . Magnetic reconnection occurs when two oppositely directed magnetic fields come very close to each other. It leads to the rearrangement of the magnetic field lines by breaking and cross-connecting to each other which results in a burst of energy. Numerical simulations using magnetohydrodynamic codes have also shown the evolution of the magnetic field and the formation of loops of magnetic field lines due to magnetic reconnection Soumen_2022 in cosmic string wakes.
Previously it has been predicted that magnetic reconnection can lead to a Gamma Ray Burst (GRB) Granot . As per the current literature, the different models of magnetic reconnection predict that the burst of energy due to magnetic reconnection is accompanied by the generation of shock waves yamada_review . In this work, we show that the collision of these shocks in the cosmic string can lead to a GRB. GRB’s are the bursts of energies observed by various telescopes. Their extreme luminosity and high energy output make them crucial for studying the early universe Paczynski_1998 . In this work, we obtain the possible GRB light pulse which will be generated from the magnetic reconnection in the cosmic string wake. As we will explain in detail, there is the possibility of multiple reconnection points occuring on the cosmic string wake and so we also obtain the synthetic GRB light curve based on our model of the light pulse.
@@
The GRB’s are generally classified based on their luminosity and the time duration over which they occur piran . These are two of the basic features of any GRB. The short duration GRB’s last for less than two seconds while the long duration GRB’s usually last for more than two seconds. We find the GRB light curves which indicate that short duration GRB’s can be generated from magnetic reconnection in cosmic string wakes. We also find that the possible time durations and the fluence calculated in our model is within the range of the short duration GRB’s identified by satellites like BATSE (Burst And Transient Source Experiment) batse . Overall, we establish that magnetic reconnection in cosmic string wakes may provide a new way of identifying these exotic objects using both electromagnetic signals and signatures from density inhomogeneities.
II Magnetic Reconnection in String Wakes
In a recent work, we have shown that the small scale inhomogeneities which are perpendicular to the shock fronts in the cosmic string wakes generate magnetic fields in the wake due to the Biermann battery mechanism Sovan_2020 . Due to the symmetry of the wake, the fields generated are oppositely directed on the two sides of the wake. As the width of the wake is small, and the background particles stream into the wake overdensity, they push the magnetic lines together and magnetic reconnection occurs. A basic calculation based on the Sweet–Parker (SP) model Parker_1957 shows that a large amount of energy is released due to the magnetic reconnection in the cosmic string wake. The SP model approximated the problem as a two dimensional problem with a plasma of constant resistivity. The energy though large was released over a substantial period of time. Other models increased the reconnection rate by using a variable resistivity che . According to the SP model, the outflow velocity is dependent on the reconnection velocity. This velocity is given by, $V_{r}=\sqrt{{C_{H}\eta c}\over{4\pi l}}$. Here, $C_{H}=\frac{B_{0}}{\sqrt{4\pi\rho}}$ is the velocity with which the magnetic lines of force move towards each other, $\eta$ is the resistivity and $l$ is the length of the diffusion region, which is the region over which the magnetic reconnection takes place. The velocity $C_{H}$ is referred to as the hydromagnetic velocity. Shocks are generated at the points of reconnection and these shocks move in an outward direction away from the point of reconnection. The shocks are relativistic shocks and have different velocities and gamma factors. Since these shocks have different velocities, the shocks will collide with one another. In Fig.[1], we show the region in the cosmic string wake where the magnetic field lines come together and points of magnetic reconnection are formed.
Figure 1: Points of magnetic reconnection being formed in the cosmic string wake. Here X marks the position of the cosmic string.
@@ II Magnetic Reconnection in String Wakes
At any given time $t_{i}$, the cosmic string wake has dimensions given by branden1 , $c_{1}t_{i}\times t_{i}v_{s}\gamma_{s}\times\delta\theta t_{i}v_{s}\gamma_{s}$ where, $c_{1}$ is of order $1$, $v_{s}$ is velocity of the moving string, $\gamma_{s}$ is the Lorentz factor associated with the velocity of the string. Since the string usually moves at relativistic velocities, the average velocity is given by $<v_{s}\gamma_{s}>$ $\sim 0.5$ dacunha . The $\delta\theta$ in the dimensions is the opening angle of the cosmic string wake. It is equal to the deficit angle of the space time around the cosmic string and is given by $\delta\theta\sim 4\pi G\mu$. The value of $G\mu$ is given by the symmetry breaking scale at which the cosmic string is formed vilenkin . Since this is very small (of the order of $10^{-7}$ or less), the width and the height of the wake is also very small compared to the length of the wake. The oppositely directed magnetic fields are along the $z-$direction so the relevant length would be along the $z-$direction. Usually a quasi 2-dimensional model is used to understand magnetic reconnection, the dimensions of the cosmic string wake also indicate that the two oppositely directed magnetic field are separated by a very small lengthscale.
As the string moves, the fluid motion due to the cosmic string metric pushes the fluid towards the center of the wake. The magnetic field lines also feel this hydrostatic pressure and are therefore pushed towards the center of the wake. As the magnetic field lines come very close to each other, points of magnetic reconnection occur. As the string moves forward, multiple such points of magnetic reconnection can happen. The shocks produced by these multiple points of reconnection will move outward. These shocks will have different velocities and the collision of these relativistic shocks will result in the generation of a GRB.
We consider a model where a slow shock is followed by a faster shock so that we can calculate the total energy of the plasma at the time the two shocks collide and merge with each other. Though there may be multiple shocks, for simplicity, we consider the case of only two shocks colliding with one another. We calculate an approximate value of the energy that is generated due to the collision of the shocks and the burst duration. Our basic assumption is that the shocks have different relativistic velocities and hence different $\gamma$ values. We have assumed that the first shock is the slower shock and has a relativistic velocity given by $\gamma_{1}v_{1}$ and the second shock is the faster shock having a relativistic velocity $\gamma_{2}v_{2}$. After the collision, the two shocks merge together and move as one with the merge density $\rho_{m}$ and merge Lorentz factor $\gamma_{m}$. The total kinetic energy released by the merger of the shocks is given by,
$K.E=\frac{1}{2}\rho_{m}\gamma_{m}^{2}c^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l}.$ (1)
For different models of reconnection, the length of the diffusion region $l$ is what makes a difference in the rate of reconnection and the energy emitted Kulsrud ; yamada_review . In the SP model, it value is over the whole region where reconnection occurs whereas in the other models it is scaled due to the emergent shocks which increases the overall reconnection rate. The time duration of the energy released is $\Delta t=\frac{r_{col}}{2\gamma_{1}^{2}c}$ beniamani . Here, $r_{col}$ denotes the radius over which the collision takes place. This too is model dependent.

### v2

@@ Abstract
Magnetic reconnection in magnetized wakes of cosmic strings results in the release of a large amount of energy. This energy is released in a short period of time. In this work, we show that this sudden release of energy can result in a Gamma Ray Burst (GRB) of short duration. The magnetic reconnection occurs at several points of the cosmic string wake. The emerging shocks from these points have different velocities. These shocks will collide with each other and give rise to short bursts of energy. The emitted pulse of energy depends on the background magnetic field and the timescale associated with the magnetic reconnection in the cosmic string wake. We also obtain the synthetic lightcurve that occurs from multiple collisions of the shock waves that are emitted from the multiple points of magnetic reconnection in the wake region. Finally, we fit the current experimental data of short GRB which are in the same energy range as the ones predicted from our model.
@@ I Introduction
Symmetry breaking phase transitions in the early universe give rise to topological defects. One of the topological defects that is predicted by these symmetry breaking phase transitions is the cosmic string defect. The cosmic string defect is a one dimensional defect generated due to a $U(1)$ symmetry breaking phase transition. Once formed, these strings will move through the early universe plasma. Depending on the exact nature of the phase transition, long strings or loops of cosmic strings are formed. Some of these defects are unstable and decay with the evolution of the universe, others are stable and evolve as a giant network of cosmic strings spanning the universe cosmicstrings . Theoreticians have predicted multiple observational signatures of these cosmic strings. Primary amongst these were the gravitational lensing of objects due to the conical space time of the cosmic string lensing . Other than the gravitational lensing, as the cosmic string moves through the plasma, it gives rise to wakes of overdensity behind it wakes . These wakes will lead to a distinct wedge shaped patterns in the $21$ cm redshift survey maps branden1 ; dacunha . These wakes also give rise to temperature fluctuations which have been used for searches for these exotic objects by the Planck collaboration Planck_2016 . Constraints on the deficit angle of the cosmic string has been studied from the radio maps of CMB anisotropy, provided by the space mission Planck for various frequencies Sazhina_2014 . All these observational signatures are currently from density inhomogeneities generated as the cosmic string moves through the plasma. Recently, there have been some signatures of Gravitational waves generated by these cosmic string networks Blasi_2021 .
@@
The wakes of these cosmic strings are long and narrow. The Biermann battery mechanism ohira produces magnetic fields that are oppositely directed along the two sides of the wake. In the next section, we will give a brief discussion on the generation of the oppositely directed magnetic fields in the cosmic string wake and the neutral line between the fields where the points of magnetic reconnection occur. Detailed discussion on these phenomena can be found in our previous studies reported in ref.Sovan_2020 and ref.Dilip_2023 .
@@
Magnetic reconnection occurs when two oppositely directed magnetic fields come very close to each other. It leads to the rearrangement of the magnetic field lines by breaking and cross-connecting to each other which results in a burst of energy. Numerical simulations using magnetohydrodynamic codes have also shown the evolution of the magnetic field and the formation of loops of magnetic field lines due to magnetic reconnection Soumen_2022 in cosmic string wakes. We had in our previous work predicted that this magnetic reconnection can lead to a Gamma Ray Burst (GRB) Dilip_2023 . GRB’s are the bursts of energies observed by various telescopes. Their extreme luminosity and high energy output make them crucial for studying the early universe Paczynski_1998 .
Previously GRBs have been associated with superconducting cosmic strings Babul_1987 , where the electromagnetic radiation is powered by the electric currents flowing through the string cores. Cusps of superconducting strings also emit powerful beamed pulse of electromagnetic radiation berezinsky which can be identified with GRBs. The only possibility of GRBs from non-superconducting cosmic strings was our prediction of a GRB from magnetic reconnection in the magnetized wake of a cosmic string Dilip_2023 . Magnetic reconnection has been considered a progenitor for a GRB in many other scenarios Granot . Depending on the different models of magnetic reconnection, shocks are usually generated at the points of reconnection yamada_review . In our previous work, we did not consider the details of the magnetic reconnection in cosmic string wakes, therefore we could give only the total energy or fluence of the GRB that is generated. In this work, we look at the points of magnetic reconnection in cosmic string wakes in greater detail and obtain a detailed light curve of a possible GRB from the wake of a magnetized cosmic string. In the summary, we will also discuss how the GRBs from the cusps of superconducting strings will be different from the ones in the wake of the cosmic string.
We have shown in a previous work which we discuss briefly in the next section that magnetic reconnection can occur in the wakes of cosmic strings. Due to the structure of the wake, multiple points of magnetic reconnection will occur. As the magnetic lines rearrange themselves, shocks can occur in the outflow. Slow mode shocks have been predicted in the Petschek model Petschek and fast mode shocks have been predicted for models with fast reconnection. In a cosmic string wake, these reconnection points are close to each other and therefore the shocks emanating from the points of magnetic reconnection will collide with one another and merge to form newer shocks. The merging of the shocks will lead to a sudden increase of the total energy in the plasma. This will cause the acceleration of the particles in the plasma. This sudden relativistic acceleration will lead to the radiation of non-thermal energy which can be identified as a GRB. We calculate the energy released during this process and obtain the shape of the pulse of a GRB light curve based on the parameters of the cosmic string wake. Since multiple such pulses will be generated, we also consider the case of large number of pulses with different velocities coming together to form the GRB light curve.
@@
The GRB’s are generally classified based on their luminosity and the time duration over which they occur piran . These are two of the basic features of any GRB. The short duration GRBs last for less than two seconds while the long duration GRBs usually last for more than two seconds. The GRB light curves we obtain are for high-redshifts and are of short-duration. We also find that the possible time durations and the fluence calculated in our model is within the range of the high-redshift, short duration GRBs identified by the satellites like FERMI and SWIFT. Overall, we establish that magnetic reconnection in cosmic string wakes may provide a new way of identifying these exotic objects using both electromagnetic signals and signatures from density inhomogeneities.
II Magnetic Reconnection in Cosmic String Wakes
In a recent work, we have shown that the small scale inhomogeneities which are perpendicular to the shock fronts in the cosmic string wakes generate magnetic fields in the wake due to the Biermann battery mechanism Sovan_2020 . We will now briefly describe this scenario in this section. We will then proceed to discuss how mulitple points of magnetic reconnection can occur in the cosmic string wake in the next subsection.
II.1 Biermann battery mechanism in the wakes of cosmic strings
@@
We know from the literature, that at any given time $t_{i}$, the cosmic string wake has dimensions given by branden1 , $c_{1}t_{i}\times t_{i}v_{s}\gamma_{s}\times\delta\theta t_{i}v_{s}\gamma_{s}$ where, $c_{1}$ is of order $1$, $v_{s}$ is velocity of the moving string, $\gamma_{s}$ is the Lorentz factor associated with the velocity of the string. Since, the string usually moves at relativistic velocities, the average velocity is given by $v_{s}\gamma_{s}>$ $\sim 0.5$ dacunha . The $\delta\theta$ in the dimensions is the opening angle of the cosmic string wake. It is equal to the deficit angle of the space time around the cosmic string and is given by $\delta\theta\sim 4\pi G\mu$. The value of $G\mu$ is given by the symmetry breaking scale at which the cosmic string is formed vilenkin . Since this is small (of the order of $10^{-7}$ or less), thus the width of the wake is small compared to the length of the wake.
The cosmic string wakes are thus narrow with the length of the order $\sim 10^{28}$ cm at around redshift $z\sim 6.7$ and a width of the order $\sim 10^{22}$ cm. The wake itself has an overdensity of a height approximately given by $t_{i}v_{s}\gamma_{s}$, where $t_{i}$ is the time at which the wake is being observed and $v_{s}\gamma_{s}$ is the relativistic speed at which the cosmic string is moving through the plasma. The plasma itself is not stationary and the plasma particles are dragged into the wake of the string due to the conical space-time of the cosmic string metric. So, both the plasma particles as well as the cosmic string are moving and the velocity we talk about is the relative velocity of the particle in the reference frame of the cosmic string. Due to this large overdensity of the wake, a temperature gradient is formed across the wake ssanyal1 ; ssanyal2 . On the two sides of the wake we have a lower temperature and the wake itself is at a higher temperature. So there are two oppositely directed temperature gradients on both sides of the wake. The wake itself is not homogeneous with significant inhomogeneities being generated due to the different masses of the particles in the wake. We have shown in ref. Sovan_2020 that a neutrino current is generated close to the cosmic string which is oscillatory in nature. This oscillatory neutrino current drives the electrons to generate a correponding oscillating electron current. Thus, electrons and positively charged particles are distributed inhomogeneously in the wake of the cosmic string. The gradient of the density inhomogeneity is roughly perpendicular to temperature gradient on the two sides of the wake, so, a magnetic field is generated in the wake. It has also been shown previously that particles of different masses cluster at different distances from the cosmic string sornborger , this again leads to density inhomogeneities in the cosmic string wake. These inhomogeneities are along the wake and therefore perpendicular to the temperature gradient. Thus according to the Biermann mechanism, they will generate magnetic fields as the term $\nabla\rho\times\nabla T$ will be non-zero (i.e. the density gradients are not aligned with the temperature gradient).
II.2 Magnetic reconnection in the wakes
As mentioned before the temperature gradient is oppositely directed in the cosmic string wake [For figure: refer to Sovan_2020 ]. So due to the symmetry of the wake, the fields generated are oppositely directed on the two sides of the wake.
Figure 1: The schematic diagram shows a cosmic string (marked by $\odot$) on the $X$-axis, moving through the plasma. The cosmic string is perpendicular to the plane of the paper. As it moves, a wake forms behind the string, outlined by solid brown lines. The opposite magnetic fields are then generated in this wake by the Biermann battery mechanism, with their directions shown by blue dashed and red solid lines. Points of magnetic reconnection being formed in the cosmic string wake shown by small dots (magenta).
In Fig.1, we have shown that the cosmic string is moving to the right while the plasma flows to the left. The two dark brown (solid) lines depict the two sides of the cosmic string wake which is symmetric about the $X$-axis of the coordinate system considered. Since we are only looking at the XYX-Y plane, the magnetic field lines (solid red line and dashed blue line) are shown on both sides of the $X$-axis. As seen from the figure, the temperature gradient is oppositely directed on two sides of the wakes along the $Y$-axis. The density inhomogeneities are along the $X$-axis. So the magnetic fields are along the positive and negative $Z$-direction (not on the depicted plane). Thus, the magnetic field from Biermann battery mechanism is given by,
$\vec{B}~=~-\hat{i}\frac{\partial N_{e}}{\partial z}\frac{\partial T}{\partial y}~+~\hat{k}\frac{\partial N_{e}}{\partial x}\frac{\partial T}{\partial y}$~ (1)
The width of the wake overdensity is small as mentioned previously so a quasi two dimensional study is justified so the subsequent field is approximated by, $\vec{B}=-\hat{i}\frac{\partial N_{e}}{\partial z}\frac{\partial T}{\partial y}$ for the quasi- 2 dimensional picture. The diffusion region is marked with the dashed line parallel to the $X$-axis and small dots (magenta color) are put to indicate the possible points of reconnection in Fig.1.
Generally due to the frozen in condition of the magnetic field, the magnetic field lines are deformed by the plasma flow. So as the background particles in the plasma stream into the wake overdensity with relativistic velocities, they push the magnetic field lines together and magnetic reconnection occurs. The oppositely directed magnetic fields are along the $Z$-direction so the relevant length would be along the $Z$-direction. As mentioned, a quasi 2-dimensional model is used to understand magnetic reconnection, the dimensions of the cosmic string wake indicate that the two oppositely directed magnetic field are separated by a very small lengthscale. So a quasi two dimensional model can be used to understand the magnetic reconnection in the cosmic string wake.
A basic calculation based on the Sweet–Parker (SP) model Parker_1957 shows that a large amount of energy is released due to the magnetic reconnection in the cosmic string wake. The SP model approximated the problem as a quasi-two dimensional problem with a plasma of constant resistivity. The energy though large was released over a substantial period of time. In the current work, we do not use the SP model. Instead we will work with faster reconnection models such as the Petschek model. There are other models which increase the reconnection rate by using a variable resistivity che .
The Lundquist number is very high for these astrophysical plasmas. The typical length of the wake is large and the resistivity is low, so the Lundquist number for our case is obtained of the order of $\sim 10^{22}$. Such high numbers have previously been found for different scenarios of astrophysical conditions lundquistno . At such high Lundquist numbers, fast magnetic reconnection occurs often leading to the formation of mulitple plasmoids.
Figure 2: The Schematic figure represents the regions of fast magnetic reconnection and plasmoids formation. The upper part of the figure shows the generation of shocks due to fast reconnection, while the lower part depicts the multiple reconnection events that lead to the formation of plasmoids. The two regions are presented separately for clarity.
In Fig.2, we have shown the region of fast magnetic reconnection in the upper part of the figure and the mulitple plasmoids being formed in the lower part of the figure. The two parts are separate. The upper one depicts the formation of shocks due to fast magnetic reconnection, while the one below shows that mulitple reconnection which form the magnetic islands or plasmoids. In a previous numerical simulation of magnetohydrodynamics in a cosmic string wake, mulitple such islands have been seen. [See for ref.Soumen_2022 for details of the numerical simulations of magnetic reconnection in cosmic string wakes.]
III Shocks generated in the wake
Hydrodynamic shocks have previously been studied in detail in the wakes of cosmic strings sornborger . The shocks are formed as the cosmic string move with relativistic speed through the early universe plasma. The particles in the plasma streaming past the moving string generate the wake as shown in Fig.1. The dark brown solid lines in Fig.1 denote the boundaries of the wakes where hydrodynamic shocks are typically generated. However, shocks in a magnetized wake has not been studied in great detail. Some recent studies show the formation of shocks in a magnetized wakes of cosmic string Soumen_2022 . Both weak as well as strong shocks can be formed in the wake depending on the relative velocity of the plasma particles and the cosmic string. Alfven’s theorem ensures the squeezing of the magnetic field lines in the cosmic string wake due to the inward streaming of the plasma particles but Alfven’s theorem breaks down at small lengthscales and multiple X-points of magnetic reconnection are obtained in the width of the wake Bist2025 .
It is well known in the literature that shocks are generated in the outflow of the Petschek and other fast reconnection models yamada_review . These shocks will move outward and on coming in contact to large density inhomogeneities will be reflected as stronger shocks. So within the cosmic string wake we will have shock-shock collisions.
Usually the shocks generated in the Petschek model are the slow mode shocks. But faster shocks can also be generated due to other models of reconnection. There are studies of fast reconnection in astrophysical plasmas. One of the context is when magnetic reconnection occurs in turbulent MHD Eyink . In such cases, the magnetic islands or plasmoids may not be aligned with the symmetry axis of the cosmic string wakes as shown in Fig.2. Thus in the literature, there is a variety of plasma reconnection scenarios which give rise to shock structures Workman and it is possible to generate fast mode shocks in magnetic reconnection outflows. Even if there are weak shocks generated in the wake of the cosmic string, they will collide with the shock from the wake boundary and be reflected back as stronger shocks. Therefore, there will be shock-shock collisions in the cosmic string wake structure. There are many different scenarios in which the collision can take place. In this case, we do the calculation for one such scenario. There are other different possibilities that may arise but we believe the results would be generically similar to the current case. In the next section, we discuss the model of Shock -Shock collision we follow and give the detailed equations.
IV Models of Shock-Shock collisions
There are many models of shock-shock collision which have been studied both analytically as well as numerically Hartigan ; Rahaman . There have also been models of collision of plasmoid blobs meng ; yuan , However we are considering the shocks as shells emanating from the reconnection points with variable Lorentz factors. One of the models that has been used in such a scenario is the hydrodynamic shell collision model Asaf . Here, the shocks are typically modeled as hydrodynamic shells of different densities and different velocities. They may be initiated from the same point or different points. Similarly the collision between the shells may be head on or may be at an oblique angle. Shocks may merge together after collision or they may be reflected by the other shock front. Here we work on only one scenario out of all these. We consider a model where a slow shock is followed by a faster shock. The energy reaches it’s peak as the two shocks merge together and move with the combined total kinetic energy. The particles of the plasma are accelerated to extremely high velocities at the moment of the collision and then decelerated as the merged shock gradually slows down. The accelerated charged particles moving in the magnetized wake of the cosmic string will then release a considerable amount of non-thermal radiation in a short time which can be identified as a GRB.
We consider the two shocks to be emanating from the same point one after the other after a small interval of time. Both the shocks move with relativistic velocities but the first shock is assumed to be slower than the second shock. We have considered that the slower shock, has a density of $\rho_{1}$ and a relativistic velocity given by $\gamma_{1}v_{1}$. The second shock is the faster shock with a density of $\rho_{2}$, having a relativistic velocity $\gamma_{2}v_{2}$. The kinetic energy density of individual shocks from magnetic reconnection is given by Dilip_2023 ,
$(K.E.)_{i}=\dfrac{1}{2}\rho_{i}\gamma_{i}^{2}c^{2}\frac{B}{(4\pi\rho)^{1/2}\sigma l}.$ (2)
The subscript $i$ will be $1$ or $2$ depending on the shock being considered and $\rho$ is the density of the plasma. After the collision, the two shocks will merge together and move as a single shock. So, after collision we denote the merged density as $\rho_{m}$ and the merged Lorentz factor as $\gamma_{m}$. For different models of reconnection, the length of the diffusion region $l$ is what makes a difference in the rate of reconnection and the energy emitted Kulsrud ; yamada_review . In the fast reconnection models, the length is scaled due to the emergent shocks which often increases the overall reconnection rate. In the Petschek model Petschek , the length of the diffusion region $l$ is smaller than the diffusion region of Sweet-Parker model, and the width of diffusion region can be calculated from this length. The width is further scaled by the ratio of the magnetic fields.